# Towards Robust SDRTV-to-HDRTV via Dual Inverse Degradation Network

### v2

@@ Abstract
Index Terms:
Inverse Tone Mapping , Video Enhancement , Neural Network , Wavelet Transform , High Dynamic Range Television
I Introduction
@@
High dynamic range television (HDRTV) has become increasingly popular because it can more realistically reproduce real-world luminance and color information, providing people with a better video viewing experience. The main differences between SDRTV and HDRTV are dynamic range, color gamut, and bit depth. However, despite the advances in HDRTV technology, there is a lack of available HDRTV content. The comparison between HDRTV and SDRTV is shown in Table I. Therefore, the conversion of SDRTV to HDRTV is an important work as it can help to alleviate the scarcity of HDRTV content and improve the video viewing experience for consumers.
Recently, convolutional neural networks are well suited for low-level image and video enhancement. There have been a large number of specific applications, such as image super-resolution[1, 2, 3, 4, 5, 6, 7], image denoising[8], image synthesis[9] and so on. Notably, the CNN-based method has emerged to convert SDRTV to HDRTV.
[10, 11]Design different methods to extract SDRTV priors and modulate latent features with the help of prior features, which can convert SDRTV to HDRTV frame by frame. [12] proposed a frequency domain processing method to convert SDRTV into HDRTV. [13] proposed a method of sub-brightness area processing to process the bright areas and dark areas of SDRTV separately to obtain high-quality HDRTV. Nonetheless, it must be acknowledged that certain issues remain and require further investigation. Through a comprehensive analysis combining practical requirements for SDRTV to HDRTV conversion with existing technologies, three key observations were identified.
@@
Color gamut
Rec. 2020[14]
Rec. 709[15]
Brightness range
$>$ 1000 nits [16]
100 nits
Color depth
10-bit
8-bit
Contrast ratio
Higher
Lower
Dynamic range
High
Standard
Content availability Lack
Sufficient
Table I: HDRTV VS SDRTV. Compared with SDRTV, HDRTV has obvious advantages in visual quality. However, HDRTV video resources are scarce, so converting SDRTV videos to HDRTV videos is of great significance.
@@
Figure 4: Performance and speed comparison with previous methods in terms of PSNR, runtime and parameter.
II Related Work
SDRTV-to-HDRTV conversion is the reconstruction of standard dynamic range video (SDRTV) images into high dynamic range video (HDRTV). Note that SDRTV-to-HDRTV is not the same as HDR imaging, which synthesizes linear irradiation images through multiple frames of LDR images. We briefly introduce some HDR imaging methods. HDR imaging methods vary, focusing on single-exposure reconstruction, multi-exposure reconstruction, and image quality assessment. Single-exposure techniques[19, 20] imitate the LDR formation pipeline using different camera response functions, while multi-exposure methods cite9116898 merge several LDR images of varying exposures. Quality assessment[21] involves metrics like Peak Signal-to-Noise Ratio, Mean Squared Error, Structural Similarity, and HDR-VDP-2, evaluating HDR images based on criteria like structural differences and visual quality. These diverse approaches highlight the ongoing advancements in HDR imaging research.
[22, 23] first studied the problem of super-resolution and SDRTV-to-HDRTV. In these works, the input image is decomposed into a detail component for texture reconstruction and a base component for contrast enhancement. Specifically, [23] first performs the SDRTV-to-HDRTV conversion using a convolutional neural network (CNN). Then, [22] introduces modulation blocks to modulate the local intensity in a spatially varying manner to achieve adaptive local contrast enhancement. Recently, [10] has proposed a scheme for SDRTV-to-HDRTV. Inspired by the SDR/HDR formation process, [10] proposes a three-step solution pipeline that includes adaptive global color mapping, local enhancement, and highlight generation. [10] also provides a benchmark dataset called HDRTV1K for SDRTV to HDRTV conversion. [24, 11] proposes a model for joint local and global feature modulation[25, 26, 27] capable of local adaptive tuning. [24] proposes a feature mapping model and uses dynamic convolution to model feature transformations, thus completing the inverse tone mapping process more accurately.
@@ III-A Motivations
III-B Overall of Dual Inverse Degradation Network
@@
III-C Restoration: Temporal-Spatial Alignment Fusion Module
@@
III-D Auxiliary Supervision: Restoring High Quality SDRTV
@@
III-E Feature Frequency Enhancement: Better Restoration of High-Frequency Detail in HDRTV
@@
III-F Dual Modulation Inverse Tone Mapping
III-F1 Preliminary of Global Feature Modulation
@@
III-F2 Feature Modulation = Convolutional Kernel Modulation
@@
III-F3 Dual Modulated Convolution
@@ IV-A Experiment Settings
IV-B Quantitative Results
Table II: Performance Comparison of Different SDRTV-to-HDRTV Methods based on PSNR.
@@ IV-B Quantitative Results
Table III: Performance Comparison of Different SDRTV-to-HDRTV Methods based on SSIM.
@@
Results On PSNR. Our proposed SDRTV-to-HDRTV network models, DIDNet and DIDNet(Tiny), exhibit a clear advantage in PSNR compared to other methods, showcasing superior results. Detailed results are shown in Table II. Taking the average PSNR as an example, DIDNet achieves PSNR values of 35.39, 34.26, 32.89, and 31.24 at QP=27, 32, 37, and 42, respectively. Meanwhile, DIDNet(Tiny) achieves PSNR values of 35.06, 34.04, 32.70, and 31.12 at the same QP settings. This indicates that our methods consistently achieve higher quality when converting SDRTV to HDRTV in various compression-quality settings.
@@
Table IV: Performance Comparison of Different SDRTV-to-HDRTV Methods based on MS-SSIM.
@@
Table V: Performance Comparison of Different SDRTV-to-HDRTV Methods based on $\Delta E_{ITP}$.
@@
Table VI: Performance Comparison of Different SDRTV-to-HDRTV Methods based on VSI,VIFp and Harrpsi.
@@
Results On SSIM and MS-SSIM. In the context of SDRTV-to-HDRTV image transformation, our proposed network models, DIDNet and DIDNet(Tiny), have demonstrated a substantial and consistent superiority as quantified by the Structural Similarity Index (SSIM), a critical metric for evaluating image similarity. Through a meticulous analysis of SSIM values across a spectrum of Quantization Parameter (QP) settings, our models consistently outperform their contemporaries, steadfastly upholding heightened image similarity during the intricate conversion process from SDRTV to HDRTV. Notably, at the challenging QP=27 configuration, they achieve markedly elevated SSIM values, thereby establishing the robustness and efficacy of our methodologies. We present a tabulation of experimental results of SSIM evaluation metrics in Table III.
@@
Compared with previous methods, DIDNet achieves higher conversion quality at multiple compression quality factors, including the advantage of obtaining higher MS-SSIM scores at high compression quality. Table IV shows the corresponding detailed experimental results.
@@
Results on $\Delta E_{ITP}$. Our DIDNet model shows significant advantages in the $\Delta E_{ITP}$ (color difference) metric, whose lower value means higher color fidelity. Compared with other models, DIDNet can achieve smaller DeltaEITP values at various compression quality factors (QP), further demonstrating its superior performance in color fidelity. Of particular note is that DIDNet is still able to maintain low DeltaEITP values under high QP values, which highlights its robustness in color fidelity. The experimental results about $\Delta E_{ITP}$ are shown in Table V.
@@
Results On VSI,VIFp and Harrpsi. In the context of the Visual Saliency Index (VSI) evaluation, DIDnet demonstrates remarkable superiority across various Compression Quality Factors (QP). Its VSI scores consistently outperform those of its peer models, highlighting DIDnet’s adeptness in preserving regions of visual saliency within images. This underscores its outstanding performance in emphasizing the quality of crucial information within images. When evaluated using the Harrpsi (Local Similarity) metric, DIDnet also exhibits exceptional performance. It consistently achieves high Harrpsi scores across different QP values, indicating its prowess in preserving fine-grained details and texture information within image regions. Particularly, it excels in maintaining local similarity in image regions. Furthermore, about the VIFp (Visual Information Fidelity) metric, DIDnet once again demonstrates its competitive advantage. Its VIFp scores remain consistently high across various QP values, signifying its proficiency in maintaining the fidelity of visual information within images. This underscores DIDnet’s ability to closely mimic the processing mechanisms of the human visual system, further enhancing its performance in preserving image quality. Detailed experimental results are in Table VI
Inference speed experiments results. We show the inference speed of different methods in Fig.4. Compared with the previous SOTA method HyCondITM and the fusion method that concatenates the video restoration model (STDF) with HyCondITM, our proposed DIDNet has advantages in both speed and HDRTV conversion quality. In the trade-off between inference speed and transformation quality, our method reaches a new Pareto front.
IV-C Qualitative Results
@@ IV-D Ablation Study
Table VII: Ablation study on pre-video restoration network, Auxiliary loss (Aux loss), and 3D ConditionNet.
@@
Table VIII: Detailed ablation results of the wavelet attention module, we calculate the PSNR of the high-frequency part of the wavelet as well as Harrpsi of the model output results, which can evaluate the fidelity of high-frequency information.
@@
Table IX: Ablation study on dual modulated convolution and wavelet attention in terms of PSNR (dB). WA and DMC represent Wavelet Attention and Dual Modulated Convolution.
@@
Ablation of Video Restoration + Inverse Tone Mapping. To verify the effectiveness of the proposed method from components such as auxiliary loss, we connect the inverse tone mapping network with the video restoration network. Specifically, we designed STDF+HDRTVNet, STDF+HyCondITM, and STDF+HDRVDM. STDF is currently the most advanced video recovery method (more parameters). The specific experimental results are shown in tableVII. Compared with STDF+HDRTVNet, STDF+HyCondITM, and STDF+HDRVDM, our method has significant advantages, $\Delta$ PSNR $>0.25$. This demonstrates that the performance advantage of our approach comes from a decoupled learning approach rather than simple multi-frame alignment.
Ablation of Auxiliary Loss. Since high-quality HDRTV to low-quality SDRTV is a dual degradation process, the model needs to learn both quality enhancement (artifact removal) and inverse tone mapping. A single-model learning dual restoration suffers from coupled learning problems, which leads to degraded model performance. To solve this problem, we use the auxiliary loss to supervise the dual degradation learning process. We show the experimental results of drop auxiliary loss in Table VII. After dropping the auxiliary loss, the PSNR is dropped by 0.35. This proves that the intermediate auxiliary loss is crucial for learning the dual degradation process.
@@
Ablation of Wavelet Attention. To ablate the Wavelet Attention module, we add the $WA$ module on the basis of the previous module. The $WA$ module enhances the features in the frequency domain, which can reconstruct more details and finer edges, and the results are shown in Table IX. The average PSNR improved from 33.31 to 33.41. The gain of $WA$ to high-frequency information is verified by the PSNR of the high-frequency subband of the image. As shown in Table VIII, after the introduction of $WA$, the high-frequency PSNR increases from 40.03 to 40.35.
Table X: Detailed ablation of DMC module. The DMC (Dual Modulation Convolution) module has smaller color difference and higher color fidelity when the calculation amount is much less than the feature modulation.
@@
Table XI: Comparison of the cost of global feature modulation and convolutional kernel modulation. The computational cost of convolution kernel modulation is lower.
@@
Ablation of Dual Modulated Convolution. Our motivation for designing DMC is to enhance the tone mapping ability (under the condition of reducing the amount of calculation) As shown in Table XI, the cost of modulated convolution is significantly less computationally than feature modulation. We conducted ablation, and $DMC$ can improve HDRTV quality. It is reported from Table IX that after adding $DMC$, the PSNR of converted HDRTV is increased by 0.03. As shown in Table X, after introducing $DMC$, the color difference is reduced from 14.31 to 13.97, the color fidelity is improved. This verifies that the proposed $DMC$ module can perform the inverse tone mapping process quickly and accurately.
IV-E Video Consistency
Table XII: Experimental results of video consistency. Our method is designed with deformable convolution and 3D conditional convolution, so it can fully utilize the video timing information, so our method has obvious advantages in temporal continuity.
@@ IV-E Video Consistency
We calculate the MSRL-PSNR[48] metric, which can evaluate the video consistency of the proposed method. And we calculated the standard deviation of $\Delta E_{ITP}$, which can evaluate the color fidelity stability of the video sequence. [48] proposed a multi-scale continuous loss, and we calculated PSNR based on this loss to obtain MSRL-PSNR. Table XII shows the experimental results, in which our method outperforms the previous best solutions in terms of video consistency.
V Conclusion

### v3

@@ Abstract
keywords:
[…]
1 Introduction
@@
High dynamic range television (HDRTV) has become increasingly popular because it can more realistically reproduce real-world luminance and color information, providing people with a better video viewing experience. The main differences between SDRTV and HDRTV are dynamic range, color gamut, and bit depth. However, despite the advances in HDRTV technology, there is a lack of available HDRTV content. The comparison between HDRTV and SDRTV is shown in Table 1. Therefore, the conversion of SDRTV to HDRTV is an important work as it can help to alleviate the scarcity of HDRTV content and improve the video viewing experience for consumers.
Recently, convolutional neural networks are well suited for low-level image and video enhancement. There have been a large number of specific applications, such as image restoration[16],image derain[5, 34, 35], image deblur[23], image enhancement[51] and so on. Notably, the CNN-based method has emerged to convert SDRTV to HDRTV.
A large number of researchers have previously introduced deep neural network[37, 47] into the field of HDR imaging. The research goal of HDR imaging is to reconstruct irradiation images, which is different from HDRTV technology. [3, 31, 36, 53]Design different methods to extract SDRTV priors and modulate latent features with the help of prior features, which can convert SDRTV to HDRTV frame by frame. [46, 45] proposed a frequency domain processing method to convert SDRTV into HDRTV. [9] proposed a method of sub-brightness area processing to process the bright areas and dark areas of SDRTV separately to obtain high-quality HDRTV. Nonetheless, it must be acknowledged that certain issues remain and require further investigation. Through a comprehensive analysis combining practical requirements for SDRTV to HDRTV conversion with existing technologies, three key observations were identified.
@@
Color gamut Rec. 2020[55] Rec. 709[17]
Brightness range $>$ 1000 nits [42] 100 nits
Color depth 10-bit 8-bit
Contrast ratio Higher Lower
Dynamic range High Standard
Content availability Lack Sufficient
Table 1: HDRTV VS SDRTV. Compared with SDRTV, HDRTV has obvious advantages in visual quality. However, HDRTV video resources are scarce, so converting SDRTV videos to HDRTV videos is of great significance.
@@
2 Related Work
SDRTV-to-HDRTV conversion is the reconstruction of standard dynamic range video (SDRTV) images into high dynamic range video (HDRTV). Note that SDRTV-to-HDRTV is not the same as HDR imaging, which synthesizes linear irradiation images through multiple frames of LDR images. We briefly introduce some HDR imaging methods. HDR imaging methods vary, focusing on single-exposure reconstruction, multi-exposure reconstruction, and image quality assessment. Single-exposure techniques[38, 24, 43, 28, 50, 49, 10] imitate the LDR formation pipeline using different camera response functions, while multi-exposure methods[27] merge several LDR images of varying exposures. Quality assessment[4] involves metrics like Peak Signal-to-Noise Ratio, Mean Squared Error, Structural Similarity, and HDR-VDP-2, evaluating HDR images based on criteria like structural differences and visual quality. These diverse approaches highlight the ongoing advancements in HDR imaging research.
[19, 21] first studied the problem of super-resolution and SDRTV-to-HDRTV. In these works, the input image is decomposed into a detail component for texture reconstruction and a base component for contrast enhancement. Specifically, [21] first performs the SDRTV-to-HDRTV conversion using a convolutional neural network (CNN). Then, [19] introduces modulation blocks to modulate the local intensity in a spatially varying manner to achieve adaptive local contrast enhancement. Recently, [3] has proposed a scheme for SDRTV-to-HDRTV. Inspired by the SDR/HDR formation process, [3] proposes a three-step solution pipeline that includes adaptive global color mapping, local enhancement, and highlight generation. [3] also provides a benchmark dataset called HDRTV1K for SDRTV to HDRTV conversion. [11, 31, 10] proposes a model for joint local and global feature modulation[15, 44, 18] capable of local adaptive tuning. [11] proposes a feature mapping model and uses dynamic convolution to model feature transformations, thus completing the inverse tone mapping process more accurately.
@@ 3.1 Motivations
3.2 Overall of Dual Inverse Degradation Network
@@
3.3 Restoration: Temporal-Spatial Alignment Fusion Module
@@
3.4 Auxiliary Supervision: Restoring High Quality SDRTV
@@
3.5 Feature Frequency Enhancement: Better Restoration of High-Frequency Detail in HDRTV
@@
3.6 Dual Modulation Inverse Tone Mapping
3.6.1 Preliminary of Global Feature Modulation
@@
3.6.2 Feature Modulation = Convolutional Kernel Modulation
@@
3.6.3 Dual Modulated Convolution
@@ 4.1 Experiment Settings
4.2 Quantitative Results
Table 2: Performance Comparison of Different SDRTV-to-HDRTV Methods based on PSNR.
@@ 4.2 Quantitative Results
Table 3: Performance Comparison of Different SDRTV-to-HDRTV Methods based on SSIM.
@@
Results On PSNR. Our proposed SDRTV-to-HDRTV network models, DIDNet and DIDNet(Tiny), exhibit a clear advantage in PSNR compared to other methods, showcasing superior results. Detailed results are shown in Table 2. Taking the average PSNR as an example, DIDNet achieves PSNR values of 35.39, 34.26, 32.89, and 31.24 at QP=27, 32, 37, and 42, respectively. Meanwhile, DIDNet(Tiny) achieves PSNR values of 35.06, 34.04, 32.70, and 31.12 at the same QP settings. This indicates that our methods consistently achieve higher quality when converting SDRTV to HDRTV in various compression-quality settings.
@@
Table 4: Performance Comparison of Different SDRTV-to-HDRTV Methods based on MS-SSIM.
@@
Table 5: Performance Comparison of Different SDRTV-to-HDRTV Methods based on $\Delta E_{ITP}$.
@@
Table 6: Performance Comparison of Different SDRTV-to-HDRTV Methods based on VSI,VIFp and Harrpsi.
@@
Results On SSIM and MS-SSIM. In the context of SDRTV-to-HDRTV image transformation, our proposed network models, DIDNet and DIDNet(Tiny), have demonstrated a substantial and consistent superiority as quantified by the Structural Similarity Index (SSIM), a critical metric for evaluating image similarity. Through a meticulous analysis of SSIM values across a spectrum of Quantization Parameter (QP) settings, our models consistently outperform their contemporaries, steadfastly upholding heightened image similarity during the intricate conversion process from SDRTV to HDRTV. Notably, at the challenging QP=27 configuration, they achieve markedly elevated SSIM values, thereby establishing the robustness and efficacy of our methodologies. We present a tabulation of experimental results of SSIM evaluation metrics in Table 3.
@@
Compared with previous methods, DIDNet achieves higher conversion quality at multiple compression quality factors, including the advantage of obtaining higher MS-SSIM scores at high compression quality. Table 4 shows the corresponding detailed experimental results.
@@
Results on $\Delta E_{ITP}$. Our DIDNet model shows significant advantages in the $\Delta E_{ITP}$ (color difference) metric, whose lower value means higher color fidelity. Compared with other models, DIDNet can achieve smaller DeltaEITP values at various compression quality factors (QP), further demonstrating its superior performance in color fidelity. Of particular note is that DIDNet is still able to maintain low DeltaEITP values under high QP values, which highlights its robustness in color fidelity. The experimental results about $\Delta E_{ITP}$ are shown in Table 5.
@@
Results On VSI,VIFp and Harrpsi. In the context of the Visual Saliency Index (VSI) evaluation, DIDnet demonstrates remarkable superiority across various Compression Quality Factors (QP). Its VSI scores consistently outperform those of its peer models, highlighting DIDnet’s adeptness in preserving regions of visual saliency within images. This underscores its outstanding performance in emphasizing the quality of crucial information within images. When evaluated using the Harrpsi (Local Similarity) metric, DIDnet also exhibits exceptional performance. It consistently achieves high Harrpsi scores across different QP values, indicating its prowess in preserving fine-grained details and texture information within image regions. Particularly, it excels in maintaining local similarity in image regions. Furthermore, about the VIFp (Visual Information Fidelity) metric, DIDnet once again demonstrates its competitive advantage. Its VIFp scores remain consistently high across various QP values, signifying its proficiency in maintaining the fidelity of visual information within images. This underscores DIDnet’s ability to closely mimic the processing mechanisms of the human visual system, further enhancing its performance in preserving image quality. Detailed experimental results are in Table 6
4.3 Qualitative Results
@@ 4.4 Ablation Study
Table 7: Ablation study on pre-video restoration network, Auxiliary loss (Aux loss), and 3D ConditionNet.
@@
Table 8: Detailed ablation results of the wavelet attention module, we calculate the PSNR of the high-frequency part of the wavelet as well as Harrpsi of the model output results, which can evaluate the fidelity of high-frequency information.
@@
Table 9: Ablation study on dual modulated convolution and wavelet attention in terms of PSNR (dB). WA and DMC represent Wavelet Attention and Dual Modulated Convolution.
@@
Ablation of Video Restoration + Inverse Tone Mapping. To verify the effectiveness of the proposed method from components such as auxiliary loss, we connect the inverse tone mapping network with the video restoration network. Specifically, we designed STDF+HDRTVNet, STDF+
HyCondITM, and STDF+HDRVDM. STDF is currently the most advanced video recovery method (more parameters). The specific experimental results are shown in table7. Compared with STDF+HDRTVNet, STDF+HyCondITM, and STDF+HDRVDM, our method has significant advantages, $\Delta$ PSNR $>0.25$. This demonstrates that the performance advantage of our approach comes from a decoupled learning approach rather than simple multi-frame alignment.
Ablation of Auxiliary Loss. Since high-quality HDRTV to low-quality SDRTV is a dual degradation process, the model needs to learn both quality enhancement (artifact removal) and inverse tone mapping. A single-model learning dual restoration suffers from coupled learning problems, which leads to degraded model performance. To solve this problem, we use the auxiliary loss to supervise the dual degradation learning process. We show the experimental results of drop auxiliary loss in Table 7. After dropping the auxiliary loss, the PSNR is dropped by 0.35. This proves that the intermediate auxiliary loss is crucial for learning the dual degradation process.
@@
Ablation of Wavelet Attention. To ablate the Wavelet Attention module, we add the $WA$ module on the basis of the previous module. The $WA$ module enhances the features in the frequency domain, which can reconstruct more details and finer edges, and the results are shown in Table 9. The average PSNR improved from 33.31 to 33.41. The gain of $WA$ to high-frequency information is verified by the PSNR of the high-frequency subband of the image. As shown in Table 8, after the introduction of $WA$, the high-frequency PSNR increases from 40.03 to 40.35.
Table 10: Detailed ablation of DMC module. The DMC (Dual Modulation Convolution) module has smaller color difference and higher color fidelity when the calculation amount is much less than the feature modulation.
@@
Table 11: Comparison of the cost of global feature modulation and convolutional kernel modulation. The computational cost of convolution kernel modulation is lower.
@@
Ablation of Dual Modulated Convolution. Our motivation for designing DMC is to enhance the tone mapping ability (under the condition of reducing the amount of calculation) As shown in Table 11, the cost of modulated convolution is significantly less computationally than feature modulation. We conducted ablation, and $DMC$ can improve HDRTV quality. It is reported from Table 9 that after adding $DMC$, the PSNR of converted HDRTV is increased by 0.03. As shown in Table 10, after introducing $DMC$, the color difference is reduced from 14.31 to 13.97, the color fidelity is improved. This verifies that the proposed $DMC$ module can perform the inverse tone mapping process quickly and accurately.
4.5 Video Consistency
Table 12: Experimental results of video consistency. Our method is designed with deformable convolution and 3D conditional convolution, so it can fully utilize the video timing information, so our method has obvious advantages in temporal continuity.
@@ 4.5 Video Consistency
We calculate the MSRL-PSNR[7] metric, which can evaluate the video consistency of the proposed method. And we calculated the standard deviation of $\Delta E_{ITP}$, which can evaluate the color fidelity stability of the video sequence. [7] proposed a multi-scale continuous loss, and we calculated PSNR based on this loss to obtain MSRL-PSNR. Table 12 shows the experimental results, in which our method outperforms the previous best solutions in terms of video consistency.
5 Conclusion